OpenStudy (anonymous):

Solve for K. 2/7(K+5/8)=16/7

OpenStudy (anonymous):

HeY!

OpenStudy (anonymous):

hey do you know the answer? I've been stuck for a while.

OpenStudy (anonymous):

Is the question\[\frac{ 2 }{ 7 }\left( k+\frac{ 5 }{ 8 } \right) = \frac{ 16 }{ 7 }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

OK. The first thing I would do is multiply both sides by 7. This will get rid of the 7ths fractions. Can you do that?

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

2(k+5/8)=16

OpenStudy (anonymous):

Very good. Now divide both sides by 2 to get rid of the 2 on the left hand side. What do you get?

OpenStudy (anonymous):

(k+5/8)=8

OpenStudy (anonymous):

Excellent. You're good at this. Now, solving for k is straight forward. Subtract 5/8 from both sides. What do you get?

OpenStudy (anonymous):

Thankyou so much for the help. I finished the rest of the question. Thankyou